OpenStudy (anonymous):

multiply 9/14 and 2/15

OpenStudy (akashdeepdeb):

\[\frac{ 9 }{ 14 } * \frac{ 2 }{ 15 } = Now~~Find~~Common~~Terms~~In~~Numerator~~and~~denominator\] \[\frac{ 3*3 }{ 2*7 } * \frac{ 2 }{ 3*5 }\] [Now cancel 3 and 2] \[\frac{ 3 }{ 7 } * \frac{ 1 }{ 5 }\] = \[\frac{3}{35}\] Understood? :)

OpenStudy (anonymous):

2/35

OpenStudy (anonymous):

understood, thanks (: